### v1

# LLaVA-3D: A Simple yet Effective Pathway to Empowering LMMs with 3D-awareness

Abstract
Recent advancements in Large Multimodal Models (LMMs) have greatly enhanced their proficiency in 2D visual understanding tasks, enabling them to effectively process and understand images and videos. However, the development of LMMs with 3D-awareness for 3D scene understanding has been hindered by the lack of large-scale 3D vision-language datasets and powerful 3D encoders. In this paper, we introduce a simple yet effective framework called LLaVA-3D. Leveraging the strong 2D understanding priors from LLaVA, our LLaVA-3D efficiently adapts LLaVA for 3D scene understanding without compromising 2D understanding capabilities. To achieve this, we employ a simple yet effective representation, 3D Patch, which connects 2D CLIP patch features with their corresponding positions in 3D space. By integrating the 3D Patches into 2D LMMs and employing joint 2D and 3D vision-language instruction tuning, we establish a unified architecture for both 2D image understanding and 3D scene understanding. Experimental results show that LLaVA-3D converges 3.5$\times$ faster than existing 3D LMMs when trained on 3D vision-language datasets. Moreover, LLaVA-3D not only achieves state-of-the-art performance across various 3D tasks but also maintains comparable 2D image understanding and vision-language conversation capabilities with LLaVA.
Figure 1: Overview of LLaVA-3D. Block (a) shows that LLaVA-3D could perform both 2D and 3D vision-language tasks. The left block (b) shows that compared with previous 3D LMMs, our LLaVA-3D achieves state-of-the-art performance across a wide range of 3D benchmarks while maintaining a comparable performance on various 2D benchmarks compared with LLaVA-1.5. The middle block (c) demonstrates that LLaVA-3D is built on the 2D LMM: LLaVA, and leverages 3D patches to endow it with 3D spatial awareness, enabling it to perform various 3D vision-and-language tasks in the physical world. The right blocks (d) and (e) highlights the significantly faster convergence and inference speeds of LLaVA-3D compared to existing 3D LMMs.
1 Introduction
Recent advancements in Large Multimodal Models (LMMs) [2, 8, 30, 43] have significantly enhanced their ability to understand and reason over visual and language inputs, leading to remarkable performance in 2D visual tasks, such as 2D perception, understanding, and reasoning. Despite their advanced perceptual and reasoning capabilities, LMMs are primarily confined to virtual interactions through images or video, lacking the critical ability to interact with the physical world. To enable their deployment in real-world applications and to facilitate the emergence of new capabilities through physical interactions, it is imperative to equip LMMs with 3D spatial intelligence.
A key aspect of 3D spatial intelligence is the ability to perceive and understand the 3D world. Similar to how 2D LMMs align 2D visual features with language models using large-scale 2D vision-language datasets, a common approach to developing 3D LMMs [12, 20, 19] involves integrating 3D features encoded from point clouds into Large Language Models (LLMs) and training them on 3D point cloud-language datasets. However, in contrast to the abundance of large-scale 2D datasets, 3D datasets remain relatively scarce. Meanwhile, there are no powerful pre-trained 3D point cloud encoders, akin to CLIP ViT [42] in 2D, to serve as a foundational model that can provide strong 3D features to LLMs.
Since real-world embodied agents typically rely on ego-centric, multi-view images as their raw observations, we aim to build a 3D foundation model based on such inputs rather than 3D point clouds. There have been attempts [18, 17] to leverage the 2D foundation models, like CLIP, alongside LLMs to advance this goal. These methods resort to 2D object segmentation results [26] to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D scene features [23]. However, this multi-view image object segmentation and object-centric feature extraction pipeline is inherently complex and computationally intensive. In contrast, 2D LMMs [8, 30, 37, 38, 2] directly leverage CLIP’s image patch features with richer, fine-grained information for effective image understanding and visual reasoning. This naturally leads to the question: Can we directly build a 3D LMM upon the strong 2D priors from 2D LMMs, bypassing the obstacles in 3D data scale and 3D encoders?
In light of recent progress in 2D LMMs, we propose a simple yet effective framework, LLaVA-3D, which extends the well-established 2D LLaVA model to efficiently comprehend the 3D world while preserving its robust 2D multimodal perception and reasoning capabilities. The core innovation in our approach is the introduction of 3D Patch, a new 3D representation that bridges 2D features within a 3D spatial context. This representation is derived by simply augmenting 2D patch-wise features with 3D positional embeddings, eliminating the need for additional complex processing. After enhancing the 2D visual tokens to 3D patch features, we explore various pooling strategies to compress the 3D patch features across extensive input frames before being fed into LLM. Fine-tuned on the existing 3D vision-language datasets, our model converges rapidly and acquires 3D spatial understanding and grounding capabilities. Furthermore, the unified model architecture allows LLaVA-3D to retain the strong 2D understanding and reasoning abilities of LLaVA through joint instruction-tuning on 2D vision-language datasets.
Our experimental results demonstrate that LLaVA-3D achieves state-of-the-art performance on a wide range of 3D tasks and benchmarks [3, 40, 39, 41, 13, 51, 6], including 3D captioning, 3D question answering, and 3D grounding while requiring significantly less training time and fewer epochs than existing 3D LMMs. Additionally, LLaVA-3D achieves comparable capabilities in 2D image understanding, reasoning, and conversation to LLaVA through joint tuning on 2D and 3D vision-language instructions.
2 Related Work
2D LMMs.
Building on the success of recent LLMs, numerous studies [37, 38, 2, 30, 8, 35] explored LMMs that can jointly process visual and linguistic information. For example, LLaVA [37, 38] aligned 2D images with language models through an image encoder and a projection layer, while BLIP2 [30] employed a more sophisticated Q-Former architecture to guide the compression of visual features using textual cues. However, most early 2D LMMs were trained on single-image datasets, limiting their ability to tackle the complexities of multi-image understanding. Recently, there has been increasing interest in expanding LMMs to handle multi-image inputs, addressing the demands of real-world scenarios. For video LMMs [34, 31, 47, 28], multi-image input forms the basis for capturing temporal or action-related dynamics across sequences of video frames. On the other hand, multi-view images of the 3D scene can implicitly reveal 3D spatial relationships and other abstract relations in the environment. Recent works  [41, 36] explored whether 2D LMMs [1, 44] can leverage multi-view images to perform spatial understanding. However, these methods primarily relied on implicit learning from the data, without directly modeling the 3D world. In contrast, our LLaVA-3D explicitly models the 3D world from multi-view images, enabling advanced 3D spatial understanding and grounding capabilities.
Injecting 3D into LLMs.
As 2D LMMs achieved substantial progress in visual perception, similar efforts have been made in the 3D domain. For 3D scene-level understanding, recent works explored ways to integrate 3D inputs such as point clouds [19, 12, 20] or multi-view images [18, 17] into LLMs to enable advanced 3D scene understanding and reasoning. An important distinction among these methods is how they construct the 3D scene representation. LL3DA [12] directly used a scene-level 3D point cloud encoder to extract the 3D scene representation. LEO [20] and Chat3D v2 [19] first segmented 3D objects from the scene point cloud using the off-the-shelf 3D instance segmentation model, and then independently extracted 3D object features with object-level 3D encoders to represent the 3D scene. On the other hand, starting from multi-view images, 3D-LLM [18] and Scene-LLM [17] resorted to manually crafted 2D object segmentation to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D point representation. Unlike these approaches, our LLaVA-3D directly builds on the well-trained 2D LMM with multi-view images as input. Utilizing the 3D position embeddings, it brings the the 2D patches within a 3D spatial context to construct 3D Patches. This 3D representation enables quick adaption of LLaVA for 3D scene understanding while preserving its strong 2D image understanding ability.
3 Method
Figure 2: LLaVA-3D Architecture. Based on LLaVA, we directly add the corresponding 3D position embeddings to 2D patch visual tokens of multi-view images to construct the 3D Patches, then the 3D Patches will undergo 3D pooling and be sent into the projection layer of LLaVA to map into the LLM space and align with the LLM using 3D-visual-language data.
Previous 2D LMMs typically consist of a visual encoder to extract 2D image features, which are then aligned with the LLM via the projection layer for joint visual and language reasoning tasks. In this section, we introduce how to bridge the 2D image features within 3D spatial context to construct 3D patches (Sec. 3.13.2), and then demonstrate the 3D-aware pooling strategies to compress the 3D patches (Sec. 3.2) and finally present the 3D-aware position encoding and decoding process (Sec. 3.4), as illustrated in Fig. 2.
3.1 Preliminary
We choose LLaVA [38], a 2D LMM, to explore building a 3D LMM based on it. LLaVA uses the pre-trained CLIP encoder to extract the 2D patch features $X_{v}\in\mathbb{R}^{c\times w\times h}$ from the input image $X\in\mathbb{R}^{3\times W\times H}$, and then align the 2D patch features $X_{v}$ into with LLM space with the projection layer. A simple multi-view image adaptation [29] for LLaVA could be extracting multi-view 2D Patch features from multi-view images $X^{\prime}_{v}\in\mathbb{R}^{V\times c\times w\times h}$ and sequentially feeding them into LLM for multi-image understanding. To empower LLaVA with 3D awareness, we introduce 3D Patch, a novel 3D representation that integrates 3D spatial information into 2D patch features.
3.2 3D Patch
Our 3D Patch representations are built upon the 2D patch features $X^{\prime}_{v}$ extracted from multi-view images with CLIP visual encoder to leverage the strong visual-semantic alignment. To construct the 3D Patches, we inject the 3D position information into the aforementioned 2d patches so that the 3D Patches can explicitly model 3D spatial information while preserving the semantic information from 2D patches. As illustrated in left block of Fig. 2, given the multi-view 2D patch features $X^{\prime}_{v}\in\mathbb{R}^{V\times c\times w\times h}$, we obtain their 3D positions $P\in\mathbb{R}^{V\times 3\times w\times h}$ in the 3D world, using nearest neighbor depth and known camera intrinsic and extrinsic parameters. The 3D positions $P$ are then encoded into 3D position embeddings $P^{\prime}\in\mathbb{R}^{V\times w\times h\times d}$ through a learnable two-layer MLP, which are subsequently added to the 2D patch visual tokens, resulting in the 3D patches $X^{\prime}_{3D}\in\mathbb{R}^{V\times w\times h\times d}$:
$X^{\prime}_{3D}=X^{\prime}_{v}+\text{MLP}(P^{\prime})$ (1)
3.3 3D Patch Pooling
While the 3D Patches equip 2D patches with 3D spatial awareness, the number of 3D patches scales directly with the number of input images. Capturing a full 3D scene often necessitates a large set of images, which significantly increases the computational overhead for large language models (LLMs). To address this, we introduce a 3D-aware pooling mechanism to reduce the number of 3D patches, as illustrated in the middle block of Fig. 2.
In the 2D image or video domain, pooling is commonly applied along the 2D spatial or temporal dimensions to compress the number of tokens and extract essential semantic information. However, for 3D scene understanding, we must pool the 3D patches based on their 3D locations to ensure these features can cover and preserve the entire scene’s structure as completely as possible. We explore two parameter-free pooling strategies to achieve this:
Voxelization Pooling.
Voxelization discretizes the 3D space into a volumetric grid. Within each occupied voxel, the 3D patches undergo average pooling, resulting in updated voxel visual tokens. Only the visual tokens from occupied voxels are passed to the LLM, with the number of tokens varying across different 3D scenes. While the number of 3D patches scales with the number of images, the number of voxel patches is only determined by the partition of voxel grids. We can easily balance the number of visual tokens and the preservation of fine-grained scene features by controlling the voxel size.
FPS Pooling.
Farthest Point Sampling (FPS) is a widely used sampling strategy to select a representative subset of points from a larger set of points cloud. We apply FPS to sample 3D patches from multi-view images to a fixed number of tokens, ensuring that the sampled tokens represent the entire scene structure. While fixing the number of tokens helps the LLM efficiently process visual information, it may also result in loss of scene information.
Compared to FPS Pooling, Voxelization Pooling offers equivalent efficiency in visual token compression while preserving more detailed scene information. Furthermore, the explicit construction of voxel grids can better handle dynamic 3D scene updates, whereas FPS Pooling excels at preserving the overall 3D scene structure. We conduct quantitative experiments in Sec. 5.7 to thoroughly assess the effectiveness of these pooling strategies.
3.4 3D-aware Position Encoding & Decoding
In the previous sections, we detailed the construction of the 3D scene representation from multi-view images, establishing the foundation for further interaction with the 3D scene. Building on this, the LLM could process multi-modal inputs such as the 3D scene, language instructions, and 3D coordinate cues to generate outputs such as language responses and 3D bounding boxes, as illustrated in the right block of Fig. 2. In this section, we introduce how the model is equipped to interpret 3D coordinate information from inputs and subsequently output precise 3D bounding boxes when specific location-related task requirements are needed.
Encoding of 3D Coordinate Input.
In scenarios such as 3D dense object captioning or object-centric question answering, the language instruction contains 3D coordinates. To handle such tasks, we introduce the 3D Coordinate Token to allow the model to integrate the provided coordinates as context into its reasoning processes. Specifically, we obtain the 3D coordinate token by feeding the 3D coordinates through a two-layer MLP. These 3D coordinate tokens are fed into LLM together with 3D Patch tokens and text tokens, enabling 3D-aware perception and reasoning.
Decoding of 3D Bounding Box Output.
The integration of the 3D coordinate token empowers the model to process 3D coordinate information from input instructions effectively. However, experiments show that directly outputting 3D bounding boxes is quite challenging for LLM, and empirically, the performance is poor. To overcome this, we leverage an approach similar to previous method  [51], introducing a specialized location token that guides the grounding module to generate accurate 3D bounding boxes. Specifically, the LLM predicts a special location token to represent a 3D box prediction when the task necessitates 3D bounding box outputs. We then derive the last layer embedding of this location token and send it into the grounding module. The grounding module utilizes the location token embedding and the 3D scene feature to predict the 3D box coordinates of the target object. This process facilitates the precise generation of 3D bounding box outputs. More details can be found in the appendix.
4 Training
To leverage the 2D priors from established 2D LMMs, we train our LLaVA-3D model based on the pre-trained LLaVA-1.5. Considering the scarcity of 3D scene-language data, our training procedure comprises two stages, each focusing on different training targets of the model.
Stage 1: 3D Patch Language Alignment.
During the first training stage, we use the region-level and scene-level caption data that describe spatial relationships among 3D objects to align the 3D patches with the LLM for enhanced 3D spatial comprehension. At this stage, the input multi-view images used are selected from sequences that correspond to specific regions or entire scenes. We freeze the vision encoder and LLM parameters, and only train the projection layer and 3D position embedding layer, encouraging efficient alignment between 3D patch features and text space. Since 3D patches are derived from CLIP features augmented with 3D positional information, the alignment between 3D Patch and LLM converges rapidly.
Figure 3: LLaVA-3D-Instruct-1M. The hybrid 2D and 3D Dataset Collection. Left: Distribution of data across categories, with the outer circle representing all categories and the inner circle illustrating data subset distribution. Right: Detailed dataset quantities.
Stage 2: Task Instruction Tuning.
During the instruction-tuning stage, LLaVA-3D is optimized to respond to complex 3D V&L tasks and maintain its inherent 2D image reasoning and instruction-following capabilities. To facilitate this capability, we collect the LLaVA-3D-Instruct-1M dataset, a hybrid collection of 2D and 3D data specifically tailored for instruction tuning. The overall distribution of the dataset collection is shown in Fig 3, more details about the instructional tuning datasets are listed in the appendix. The 2D Data of LLaVA-3D-Instruct-1M is derived from existing LLaVA-1.5 instruction tuning data, ensuring the preservation of 2D image comprehension and vision-language conversation abilities. When tuning with 2D data, we keep the vision encoder frozen and jointly train the projection layer and LLM. The 3D Data of LLaVA-3D-Instruct-1M, on the other hand, comprises data from diverse 3D QA, 3D dense captioning, and 3D grounding tasks. During the 3D data instruction tuning, the 3D position embedding layer will be added to jointly train with the other modules. Additionally, for tasks where the instruction contains 3D coordinate information or requires 3D bounding box outputs, the corresponding encoding and decoding modules will be trained together. During instruction tuning, the 3D data pathway includes the 3D position embeddings and 3D patches, while the 2D data pathway is the original LLaVA. All modules except for the 3D position embeddings to construct 3D patches are shared across 2D and 3D data. This training setup ensures that LLaVA-3D is capable of processing both 2D and 3D visual tokens effectively, and is adaptive to various task formulations and requirements.
5 Experiments
Table 1: Quantitative comparison with SOTA models on various 3D QA tasks. “C” stands for “CIDEr”, “B-4” for “BLEU-4”, “M” for “METEOR”, “R” for “ROUGE”, “Sim” for sentence similarity, and “EM@1” for top-1 exact match. Gray indicates evaluation results with refined exact-match protocol.
ScanQA (val) SQA3D (test)
C B-4 M R EM@1 EM@1
Task-specific models
Scan2Cap [13] - - - - -   41.0
ScanRefer+MCAN [49] 55.4 7.9 11.5 30.0 18.6 -
ClipBERT [27] - - - - - 43.3
ScanQA [3] 64.9 10.1 13.1 33.3 21.1 47.2
3D-VisTA [52] 69.6 10.4 13.9 35.7 22.4 48.5
Task-specific fine-tuned 3D LMMs
3D-LLM (FlanT5) [18] 69.4 12.0 14.5 35.7 20.5
LL3DA [37] 76.8 13.5 15.9 37.3 -
Chat-3D v2 [19] 87.6 14.0 - - - 54.7
LEO [20] 101.4 13.2 20.0 49.2 24.5 (47.6) 50.0 (52.4)
Scene-LLM  [17] 80 12.0 16.6 40.0 27.2 54.2
Zero-shot 2D LMMs
VideoChat2 [32] 49.2 9.6 9.5 28.2 19.2 37.3
LLaVA-NeXT-Video [28] 46.2 9.8 9.1 27.8 18.7 34.2
GPT-4V 59.6 - 13.5 33.4 - -
Gemini 68.3 - 11.3 35.4 - -
Claude 57.7 - 10.0 29.3 - -
LLaVA-3D 91.7 14.5 20.7 50.1 27.0 (45.0) 55.6 (57.6)
Table 2: Quantitative comparison on MMScan QA benchmark. “S.-BERT", “B-1”, “B-4”, “R.-L.”, “MET.” represents “Sentence-BERT", “BLEU-1”, “BLEU-4”, “ROUGE-L”, “METEOR”, respectively. Here, we report the top-1 exact match with (the refined exact-match protocol results) for “EM@1”.
Methods Setting Overall Single-target Inter-target Advanced Data-driven Metrics Traditional Metrics
ST-attr ST-space OO-attr OO-space OR SimCSE S.-BERT B-1. B-4. R.-L MET. EM@1
3D-LLM [18] Zero-Shot 28.6 37.8 18.8 13.7 26.3 15.4 20.8 40.4 40.3 13.4 1.5 17.3 6.0 6.2 (19.6)
Chat3D-v2 [19] 27.9 38.1 18.3 9.3 22.4 13.5 25.4 45.4 46.3 18.0 3.0 22.9 7.5 10.2 (19.6)
LL3DA [12] 15.8 15.5 14.7 14.2 25.2 4.3 6.4 40.7 43.6 5.4 2.1 16.4 4.4 8.3 (19.4)
LEO [20] 22.2 28.9 17.6 18.1 20.4 15.0 16.3 40.4 41.0 11.0 0.7 17.1 4.9 9.6 (18.7)
LL3DA [12] Fine-tuning 38.5 40.4 46.2 14.7 47.1 26.4 7.1 65.3 67.0 26.4 8.5 44.3 14.7 30.2 (37.6)
LEO [20] 47.8 55.5 49.5 36.1 45.6 32.1 38.4 71.2 72.2 32.0 12.5 52.1 17.7 36.6 (44.5)
LLaVA-3D Generalist 52.3 61.2 54.4 28.7 61.2 44.5 43.6 74.6 76.3 38.7 13.1 55.5 19.5 45.2 (51.4)
Table 3: Quantitative comparison with SOTA models on OpenEQA benchmark
Models Frame Accuracy
LLaMA2 [45] 0 28.3
GPT-4 [1] 0 33.5
Claude3 20 36.3
Gemini-Pro [44] 15 44.9
GPT-4V [1] 15 54.6
GPT-4V [1] 50 55.3
Human Full 86.8
LLaVA-3D 20 51.2
In this section, we conduct extensive evaluations to examine the capabilities of LLaVA-3D. To begin with, we introduce the implementation details (Sec. 5.1). Then, we compare our model’s 3D scene understanding (Sec. 5.25.35.4)and 2D image understanding (Sec. 5.5) capability with previous methods. Finally, we conduct a thorough analysis to validate the effectiveness of the components and designs of our LLaVA-3D (Sec. 5.65.7).
5.1 Implementation Details
LLaVA-3D is built upon the LLaVA-1.5-7B, utilizing their pre-trained weights from the HuggingFace library. For 3D tasks, we add the 3D position embeddings to the 2D patch visual tokens, and utilize the voxelization pooling strategy to reduce 3D patch number before passing the input visual tokens to the projection layer and LLM. The number of views $V$ is set to be 20 and voxel size is set to 0.2m. Due to the LLM context length limitation, the maximum number of 3D patch tokens after 3D-aware pooling is set to 3096. For 2D tasks, LLaVA-3D functions the same as LLaVA. All experiments are conducted on 8 × 80G A100 GPUs. We train our model for 1 epoch with a learning rate of 1e-3 and a batch size of 32 in stage 1, and fine-tune on the collected LLaVA-3D-Instruct-1M dataset, with a learning rate of 2e-5 and a batch size of 16 in stage 2.
5.2 Evaluation on 3D Question Answering
3D Question Answering requires a model to generate responses to the natural language queries questioning towards a 3D scene. In this section, we validate LLaVA-3D performance on various 3D question answering benchmarks: ScanQA [3], SQA3D [40], MMScan QA [39], and OpenEQA [41].
Spatial Understanding with ScanQA and SQA3D.
ScanQA and SQA3D are both built on ScanNet dataset. The ScanQA dataset consists of 41363 questions about 800 scenes, including 32337 unique questions. Its validation set contains 4675 questions about 71 scenes. SQA3D comprises 20.4k descriptions of 6.8k unique situations collected from 650 ScanNet scenes and 33.4k questions about these situations. Questions in ScanQA require basic recognition and 3D reasoning capabilities, and SQA3D further incorporates the situation understanding and situated reasoning into embodied 3D scene understanding. Following prior works, we adopt BLEU scores, METEOR, ROUHE-L, CIDEr and EM (“exact match”) as our evaluation metrics for ScanQA and EM for SQA3D respectively. As shown in Tab. 1, current 2D LMMs fail to achieve competitive performance with the latest 3D LMMs trained with 3D awareness, which might be attributed to the lack of explicit 3D representation. Besides, compared with task-specific fine-tuned 3D LMMs that need to be further fine-tuned on the corresponding datasets, our LLaVA-3D could perform as a generalist and achieve the SOTA performance on these benchmarks.
Coordinate Spatial Understanding with MMScan QA.
MMScan QA includes 5.2k scans from ScanNet, 3RScan, and Matterport3D, along with 116k training questions and 29k validation questions. These questions span existential inquiries, attribute understanding, and more advanced queries. Unlike ScanQA and SQA3D, some MMScan QA questions require 3D reasoning based on object coordinates, rather than relying solely on text descriptions, demanding the model capable of understanding 3D coordinates information. All data samples are classified into one of the following sub-categories: ST-attr, ST-space, OO-attr, OO-space, OR, where ST stands for Single-target, attr for attribute, OO for Object-Object, and OR for Object Region. Besides, there is a minor part of QA samples for advanced understanding and reasoning, such as situated QA related to everyday life. We present the results under GPT-4 evaluation, data-driven metrics, and traditional metrics respectively in Tab. 2. In this benchmark, the well-trained LL3DA and LEO are further fine-tuned on the full 1.2M MMScan QA training dataset. Our LLaVA-3D, trained on LLaVA-3D-Instruct-1M (which includes 440K MMScan QA training samples), achieves significantly better performance on the MMScan QA benchmark compared to the specially fine-tuned LL3DA and LEO. The results highlight the training efficiency of LLaVA-3D and its strong 3D understanding ability to serve as the generalist model.
Embodied Question Answering with OpenEQA.
OpenEQA is the first open-vocabulary benchmark designed for spatial understanding and embodied reasoning in embodied question answering, specifically in the era of foundation models. It features an automated evaluation protocol powered by LLMs, which shows strong alignment with human judgment. Our evaluations are conducted using the EM-EQA data split of OpenEQA, which includes over 1,600 high-quality, human-generated questions from diverse real-world environments. In EM-EQA, the model is required to generate a textual answer to a question based on an episode history, which typically consists of RGB images, depth information, camera poses, and intrinsic camera data. The results in Tab. 3 demonstrate that LLaVA-3D surpasses Claude3 and Gemini-Pro, and achieves comparable performance with powerful GPT-4V on this benchmark with significantly fewer model parameters.
Table 4: Quantitative Comparisons with SOTA models for 3D Dense Captioning on Scan2Cap. “C” stands for “CIDEr”, “B-4” for “BLEU4”, “M” for “METEOR”, “R” for “ROUGE”, “Sim” for sentence similarity, and “EM@1” for top-1 exact match. The n-gram metrics for Scan2Cap are governed by IoU@0.5.
Scan2Cap (Val)
C@0.5$\uparrow$ B-4@0.5$\uparrow$ M@0.5$\uparrow$ R@0.5$\uparrow$
Scan2Cap [13] 39.08 23.32 21.97 44.78
MORE [24] 40.94 22.93 21.66 44.42
SpaCap3D [46] 44.02 25.26 22.33 45.36
D3Net [7] 46.07 30.29 24.35 51.67
UniT3D [14] 46.69 27.22 21.91 45.98
3DJCG [5] 49.48 31.03 24.22 50.80
3D-VLP [25] 54.94 32.31 24.83 51.51
3D-VisTA [52] 61.60 34.10 26.80 55.00
Vote2Cap-DETR [11] 61.81 34.46 26.22 54.40
LL3DA [12] 65.19 36.79 25.97 55.06
LEO [20] 72.4 38.2 27.9 58.1
LLaVA-3D 79.21 41.12 30.21 63.41
Table 5: Quantitative comparison with SOTA models on the MMScan Captioning benchmark.
model Evaluator Type Color Shape Position Function Design Overall
LL3DA [12] GPT 10.0 26.3 40.6 38.9 67.5 21.7 33.6
LEO [20] GPT 34.9 29.7 63.0 63.7 75.0 42.7 51.3
LLaVA-3D GPT 39.9 79.2 89.1 82.2 94.1 88.0 78.8
5.3 Evaluation on 3D Dense Captioning
3D dense captioning requires the model to localize all the objects in a 3D scene and then generate descriptive sentences for each object. To evaluate our model on the 3D dense captioning tasks, we utilize the off-the-shelf segmentation model Mask3D to generate object proposals. Then we further construct the 3D coordinate tokens based on the 3D object center coordinates to guide the model to perform the task. Additionally, we utilize two types of textual instructions that prompt the model to either describe the object or describe and output the bounding box of the object. We report the performance of various methods on two 3D dense captioning benchmarks: Scan2Cap and MMScan Captioning.
Scan2Cap.
Scan2Cap requires the model to describe the object’s appearance and the spatial relations with nearby objects and output the corresponding 3D bounding box. As illustrated in Tab. 4, our method consistently outperforms the existing method on the Scan2Cap benchmark.
MMScan Captioning.
MMScan Captioning focuses on identifying common aspects of 3D objects such as Object Type, Color, Shape, Position, Function, and Design. The benchmark utilizes GPT4 to assess whether these aspects are correct in the object description. We benchmark various methods on the MMScan Captioning benchmark in Tab. 5. The results show that our method surpasses existing approaches across all metrics by a substantial margin, especially achieving a 49.5% improvement in the Color score and a 43.3% improvement in the Design score. This significant performance boost can be attributed to our 2D LMM-based architecture.
Uniquely, LLaVA-3D takes multi-view images as inputs, enabling a user-friendly feature where users can simply click on selected images to generate both 3D object captions and 3D bounding boxes. We refer to this capability as 2D Click-based 3D Dense Captioning. (as illustrated in Fig. 4).
5.4 Evaluation on 3D Visual Grounding
3D visual grounding aims to localize the target object in the 3D scene using natural language descriptions. In this section, we initially report the performance on the ScanRefer [6] benchmark. For quantitative comparisons, we include both task-specific approaches and generalist models: the state-of-the-art specialists in 3D VG and generalist models like LLM-Grounder [48], 3D-LLM [18], and Chat3D-v2 [19]. 3D-LLM [18]] uses the location tokens to discretize the 3D scene and predicts the bounding box as the location tokens that were added to the vocabulary. Chat-3D v2 [19] first detects all the objects and then identifies the object that best matches the provided description. The results in Tab. 6 demonstrate that by combining with the grounding module, our LLaVA-3D could achieve even better performance compared with the task-specific specialists model as a generalist. Such a paradigm allows LLaVA-3D to integrate with various powerful 3D grounding modules and inject the 3D understanding capability and world knowledge into the grounding process, as illustrated in Fig. 7.
Table 6: Quantitative comparison with SOTA models on various 3D VG tasks.
ScanRefer
Acc@0.25 Acc@0.5
Task-specific models
ScanRefer [6] 37.3 24.3
MVT [21] 40.8 33.3
3DVG-Trans [50] 45.9 34.5
ViL3DRel [10] 47.9 37.7
BUTD-DETR [22] 52.2 39.8
ReGround3D [51] 53.1 41.1
Task-specific fine-tuned 3D LLMs
LLM-Grounder [48] 17.1 5.3
3D-LLM [18] 30.3 -
Chat3D-v2 [19] 35.9 30.4
LLaVA-3D 54.1 42.2
Figure 4: LLaVA-3D enables the user-friendly interaction with the 3D scene across various 3D understanding and reasoning tasks. It allows the users to just click on the 2D images or the video frame to simply conduct the interactive 3D question answering and 3D dense captioning.
Table 7: Quantitative Comparisons with SOTA models on zero-shot 2D benchmarks.
[…]
MobileVLM [15] MLLaMA 2.7B 336 47.5 59.6 1289 -
InstructBLIP [16] Vicuna-7B 224 50.1 36.0 26.2
InstructBLIP [16] Vicuna-13B 224 50.7 1213 25.6
Qwen-VL [4] Qwen-7B 448 63.8 38.2
Qwen-VL-Chat [4] Qwen-7B 448 61.5 60.6 1488
Shikra [9] Vicuna-13B 224 58.8
LLaMA-VID [33] Vicuna-7B 336 65.1 1521
LLaMA-VID [33] Vicuna-13B 336 66.6 1542
LLaVA-1.5 [38] Vicuna-7B 336 58.2 65.2 1511 31.1
LLaVA-1.5 [38] Vicuna-13B 336 61.3 69.2 1531/295 36.1
LLaVA-3D Vicuna-7B 336 57.8 65.0 1502 30.9
5.5 Evaluation on 2D benchmarks
Since our model is trained on a mix of 2D and 3D datasets, we evaluate it across various 2D benchmarks to ensure it retains the 2D image understanding capabilities of the original LLaVA. As demonstrated in Tab. 7, LLaVA-3D achieves performance comparable to that of LLaVA-1.5 across several 2D image understanding benchmarks. This performance underscores the architectural superiority of our model compared to other 3D LMMs.
Table 8: Architecture Comparison on various 3D V&L Benchmark.
3D Feature Connector LLM / LMM ScanQA SQA3D Inference time
(a) (SAM + CLIP) w / PE Q-Former Vicuna-7B 21.9 49.3 900s
(b) (SAM + CLIP) w / PE Pooling + MLP Vicuna-7B 22.1 49.2 900s
(c) CLIP w / PE Pooling + MLP Vicuna-7B 23.4 51.2 0.2s
(d) CLIP w / PE Pooling + MLP LLaVA-1.5 27.0 55.6 0.2s
(e) CLIP w / PE Pooling + MLP PLLaVA 27.9 56.2 0.2s
Table 9: Effectiveness of 3D Patch Representation. Training using the instruction tuning datasets, the only difference between multi-image LLaVA and LLaVA-3D is the patch type.
Method Patch Type ScanQA SQA3D MMScan QA Scan2Cap
multi-image LLaVA 2D 24.1 52.3 32.7 29.1
LLaVA-3D 3D 27.0 (+2.9) 55.6 (+3.3) 41.5 (+8.8) 79.2 (+50.1)
5.6 Architecture Analysis
In this section, we delve deeper into the architectural benefits and efficacy of adapting a 2D large multimodal model (LMM) to a 3D LMM, as opposed to developing a 3D LMM solely from LLMs. A notable approach in the latter category is 3D-LLM, which leverages foundational 2D visual models such as SAM and CLIP for feature extraction. This method then employs 3D multi-view fusion to generate 3D point features, followed by the use of a Q-Former to condense these point features into a fixed number of tokens.
Architecture Comparison.
To ensure the fairness of the experiment as much as possible, starting from LLM, we first ablate different 3D feature encoders, and 3D-language connectors using the same instruction tuning datasets. As shown in Tab. 8, the Q-Former (a) and Pooling + MLP (b) share a similar performance on 3D V&L benchmarks. Notably, using CLIP (c) alone instead of SAM + CLIP (b) achieves better performance and significantly reduces 3D feature extraction computation time from 900s to 0.2s.
Effectiveness of 3D Patch.
Once we don’t decorate the 2D patches with corresponding 3D position embedding, LLaVA-3D will degenerate into 2D multi-image LMMs. To further ascertain the efficacy of our proposed 3D representation: 3D Patch, we conduct additional experiments across a variety of 3D question answering and 3D dense captioning benchmarks. As shown in Tab. 9, integrating 2D patches within a 3D context enhances the model’s 3D spatial understanding and reasoning capabilities, resulting in 2.9% and 3.3% improvements on the ScanQA and SQA3D benchmarks, respectively. Additionally, 3D patches are crucial for tasks that require explicit 3D world modeling, leading to significant improvements of 8.8% on the MMScan QA benchmark and an impressive 50.1% on the Scan2Cap benchmark.
Benefits from pre-trained 2D LMM.
Leveraging our foundation in 2D LMMs, our framework benefits significantly from the robust pre-training on extensive image/video-text datasets. In Tab. 8, we explore the advantages of initializing from a pre-trained 2D LMM compared to starting directly from an LLM. The experimental results consistently demonstrate that starting with a well-tuned 2D LMM substantially could enhance performance in 3D spatial understanding tasks. Besides, we observe that stronger base 2D LMM could lead to better 3D spatial understanding performance. Besides, we find that initializing from a pre-trained 2D LMM could achieve 3.5× faster training convergence speed of LLaVA-3D compared with existing 3D LMMs [12, 20]. More details could be found in our appendix.
5.7 More Analysis
To better understand the impact of different components of our LLaVA-3D, we conduct a thorough ablation study on the ScanQA and SQA3D benchmarks.
Impact of Pooling Strategy.
Given the maximum token length limitation of LLMs, we apply pooling to the 3D patch tokens extracted from multi-view images to reduce the number of tokens. However, this pooling process inevitably leads to some information loss. To understand its impact on performance, we conducted experiments to evaluate the effects of pooling. As shown in Tab. 10, the voxelization pooling strategy outperforms the FPS pooling method on 3D QA benchmarks. Reducing the voxel size in the voxelization pooling or increasing the number of 3D patch tokens obtained through FPS pooling can both enhance the model’s performance to some extent.
Table 10: Comparsion on different pooling strategies.
Pooling Stratgey Voxel Size Token Number ScanQA SQA3D
Voxelization 0.4 Dynamic 24.1 53.2
Voxelization 0.3 Dynamic 25.9 54.8
Voxelization 0.2 Dynamic 27.0 55.6
FPS - 576 25.7 54.9
FPS - 1024 26.3 55.2
Multi-View Images Sampling Strategy.
Due to the redundancy of information among multi-view images and considering computational costs, we sample $V$ views from the egocentric images of the 3D scene in our experiment. In this section, we explore the effect of different image sampling strategies during the inference stage: 1) Uniform Sampling: To achieve comprehensive coverage of the entire 3D scene, a straightforward approach is uniform sampling, which samples images evenly. 2) Text-Guided Sampling: such sampling strategy uses CLIP to select the frames related to the text instruction during inference based on the image-text CLIP similarity score. Our experiments demonstrate that these two sampling strategies share similar performance on ScanQA and SQA3D, so we choose uniform sampling for simplicity.
Number of Views.
An intuitive assumption is that sampling more views from the 3D scene will preserve more information about the 3D scene. We conduct a comparative experiment varying the number of views sampled from 3D scenes. Tab. 11 presents the Exact Match (EM) scores on ScanQA and SQA3D across different settings, revealing that the increase in EM score is marginal as the number of views increases. Additionally, the experimental results indicate that exceeding a certain number of views can degrade the model’s performance.
Table 11: Comparison on performance on 3D QA tasks under different number of multi-view images.
Number of Views Number of Tokens ScanQA SQA3D
16 9216 26.2 55.1
20 11520 27.0 55.6
24 13824 27.0 55.4
40 23040 26.7 55.2
6 Conclusion
We propose LLaVA-3D, a simple yet effective framework built upon the well-established 2D LLaVA model. LLaVA-3D extends the LLaVA’s capabilities to understand the 3D world by using 3D patches to bridge 2D features within a 3D space, enabling spatial understanding while efficiently preserving the original 2D image understanding and reasoning capability. Experimental results show that our method sets new state-of-the-art performance on a wide range of 3D tasks and benchmarks. We hope that our model will inspire new ideas for building 3D LMMs, and in the future, we plan to explore the application of LLaVA-3D in more downstream scenarios, such as robot manipulation and navigation.